A beam of light enters the flat surface of a diamond at an angle of 30 degrees from the normal. The angle of refraction in the diamond is measured to be 12 degrees from the normal. Determine the speed of light in the diamond.
Answer #1

Using Snell's law we have

n1 sin i = n2 sin r

1*sin 30 = n2* sin 12

n2 = 2.4

refractive index = speed of light in vaccuum / speed of light in material

2.4 = 3*10^8 / v

v = 1.247*10^8 m/s

so the speed of light in diamond is 1.247*10^8 m/s